# Histogram Layers for Neural “Engineered” Features

### v1

@@ Abstract
In the computer vision literature, many effective histogram-based features have been developed. These “engineered” features include local binary patterns and edge histogram descriptors among others and they have been shown to be informative features for a variety of computer vision tasks. In this paper, we explore whether these features can be learned through histogram layers embedded in a neural network and, therefore, be leveraged within deep learning frameworks. By using histogram features, local statistics of the feature maps from the convolution neural networks can be used to better represent the data. We present neural versions of local binary pattern and edge histogram descriptors that jointly improve the feature representation and perform image classification. Experiments are presented on benchmark and real-world datasets. Our code is publicly available [^1] https://github.com/Advanced-Vision-and-Learning-Lab/NEHD_NLBP.
Index Terms:
@@
I Introduction
Before the popularity of deep learning, feature engineering often (and still does) played a vital role in the fields of computer vision and machine learning. Examples of engineered features include local binary patterns [ojala1994performance] (LBP) and edge histogram descriptors (EHD) [manjunath2002introduction]. Extracting and selecting the best features (i.e., feature engineering) overall was difficult. Many of these engineered features were created through the use of histograms, but there were many parameters (such as bin centers and widths) that were difficult to tune. As an alternative to the difficult and time-consuming process of feature engineering, deep learning is used to automate the process of extracting features and performing follow-on tasks (e.g., classification, segmentation, object detection).
Convolutional neural networks (CNN) are frequently used in a variety of applications. These models extract features through a series of convolutions, aggregation functions (e.g., max and average pooling) and non-linear activation functions to improve the representation of the data for better performance. Despite these powerful expressive features learned by a CNN, these models have some downfalls. CNNs have a spatial stationary assumption; therefore, CNNs are unable to account for changes in the local statistics (i.e., statistical texture features [peeples2022histogram]) for regions of the data [taigman2014deepface, liang2021patch]. Additionally, CNNs can lead to increased computational cost in terms of training time, memory requirements, and large datasets [juefei2017local]. CNNs effectively capture structural texture features while histogram layers focus on statistical texture features [peeples2022histogram].
To mitigate these issues associated with traditional and deep learning features, alternative models have been introduced that take inspiration from both approaches. A notable method is the local binary convolutional neural network (LBCNN) [juefei2017local]. The LBCNN used a novel architecture design that lead to a less expensive model that performed comparably to standard CNNs. LBCNN was inspired by LBP; however, the LBCNN generalizes the code generation and does not account for the aggregation operation of the original LBP.
Figure 1: Features maps captured by LBP (top) and NLBP (bottom) from the FashionMNIST dataset. The proposed approach nearly reconstructs the original LBP feature. The LBP and NLBP encodings differ slightly due to the non-linearity: threshold (LBP) and ReLU (NLBP). Additionally, a square window ($3\times 3$) is used for NEHD while LBP used a radial window for the local neighborhood. There are small magnitude differences of the histogram because of the soft binning approximation used in NLBP.
@@ I Introduction
Histogram-based feature approaches (e.g., LBP and EHD) as well as CNNs are able to extract texture information [maenpaa2005texture, liu2019bow, hermann2020origins, peeples2021histogram] to represent the data. There are two types of textures: statistical and structural [peeples2021histogram, ji2022structural]. Both texture feature types are useful to represent the data to improve performance and features such as LBP use both within the feature extraction pipeline [maenpaa2005texture]. In this work, we present a generalized framework to learn histogram-based features in artificial neural networks: neural LBP (NLBP) and EHD (NEHD). The contributions of this new method are four-fold:
@@
Each neural feature can be reconstructed using convolutional and histogram layers to learn the structural and statistical texture features respectively. Our proposed neural “engineered” features can closely reconstruct the LBP and EHD features as shown in Figures 1 and 2. The LBP feature consists of both structural (i.e., encoded pixel differences) and statistical (i.e., binary code frequency) texture information. To capture the structural relationships, following [liu2017local], sparse convolutional kernels can be used to capture the difference between the center and neighboring pixels in an image. After the convolution operation is performed, an activation function (e.g., sigmoid, ReLU) can be used instead of the threshold operation. After the input image has been encoded, a histogram layer can be used to adaptively aggregate the encoded features as opposed to the fixed binning of the original LBP feature.
Figure 1 shows a comparison of LBP and the proposed NLBP. To reconstruct the LBP feature, a modified ReLU activation function can be used. The standard ReLU function sets all negative values to 00 (similar to the threshold function in LBP), but if the center pixel and neighbor pixel value are the same, the LBP threshold function will output a value of $1$. To account for this in the NLBP, if the difference between a center and neighboring pixel is 00, the ReLU function was set to output a $1$ instead of a 00 resulting in a similar encoding map as the original LBP. The histogram layer bins are initialized to the corresponding LBP code values ranging from 00 to $255$. The bin width was set to $3.75$ as this value will correspond to a narrow histogram bin.
The EHD feature also captures structural (i.e., edges) and statistical (i.e., max response frequency) texture information. The structural features can be captured using a convolutional layer with each filter corresponding to a Sobel kernel for each edge orientation, $0^{\circ}$ to $315^{\circ}$. A total of eight edge kernels were used to capture horizontal, vertical, diagonal and anti-diagonal orientations (as in the standard EHD feature). To account for direction information, the bin center values are set to the maximum response of the convolution output from the edge kernels and the input data (i.e., maximum input value $\times$ ($1+2+1$)). By changing the bin center, the binning function will now account for direction information. The anti-direction will now have the lowest response to the corresponding correct direction (i.e., $0^{\circ}$ will have lowest response for $180^{\circ}$). The same bin width value ($3.75$) was used for NEHD as in NLBP. The parameters (convolution kernels, bin centers, and bin widths) were fixed to ensure that the histogram layer can mimic the EHD feature as shown in Figure 2.
@@
II Related Work
Figure 3: Example of LBP feature value calculation. After a local neighborhood is selected (e.g., $3\times 3$), the difference between the center and neighboring pixels is computed. If the difference is greater than or equal to a chosen threshold (e.g., 00), then a value of “1” is returned for that neighbor. The next step is to encode the binary position of each neighbor (i.e. $2^{0}$, $2^{1}$, … , $2^{7}$). An elementwise multiplication and sum operation is performed between the threshold and binary position values. The output of this operation is the LBP feature for the local patch. The values from each local patch of the input image is then aggregated via a histogram to generate the output LBP histogram vector.
II-A “Engineered” Histogram Features
Histograms are used throughout computer vision and machine learning as a method to aggregate intensity and/or feature values as well as relationships between neighboring inputs (e.g., edge orientation, pixel differences). Common histogram-based features include LBP [ojala1994performance], histogram of oriented gradients (HOG) [dalal2005histograms], EHD [frigui2008detection], and gray level co-occurrence matrix (GLCM) [haralick1973textural]. For this work, we will focus on LBP and EHD.
@@ II Related Work
LBP is computed by first selecting a neighborhood of size $\mathcal{N}$ (commonly 3 $\times$ 3 or also using a radial window operation) and computing the difference between a center pixel, $x_{c}$, and each of the pixels in the defined neighborhood, $x_{i}$, in a gray scale image. The LBP feature value for a pixel in an image is computed by computing the product of the threshold, $\mathcal{T}$, of the difference between the center pixel and neighbor pixel, and two raised to the power of the index of the neighbor (Equation 1). The threshold for assigning a “0” or a “1” is usually 0 (i.e., the difference needs to be positive) as shown in Equation 2, but other variants may use different thresholds [zhou2009face, jiang2016adaptive]:
@@
LBP is robust to monotonic changes in grayscale due to the threshold function [ojala1994performance, ojala1996comparative, humeau2019texture] (i.e., if the intensity values of an image change, the binary threshold will always return either two values, “0” or “1”, despite gray level changes). There are a plethora of extensions to LBP in the literature, and several works that summarize the novelty of each LBP based approach [liu2017local, fernandez2013texture].
@@
The EHD feature records the frequency and orientation of intensity changes in an image [frigui2008detection]. A single channel input image is convolved with a set of $K$ filters (e.g., Sobel [sobel19683x3]) of size $M\times N$ to calculate the edge responses, $R\in\mathbb{R}^{M^{\prime}\times N^{\prime}\times K}$. The edge responses are grouped (“binned”) together based on their orientation such as vertical, horizontal, diagonal, anti-diagonal, and no edges (isotropic) [frigui2008detection]. The five groups can include signed [peeples2018possibilistic, peeples2019comparison] or unsigned orientations [frigui2008detection] (e.g., 0 and 180 are both horizontal but different directions). To compute the histogram of edge orientations, $\mathbf{H}\in\mathbb{R}^{K+1}$, the maximum edge responses are summed over the spatial dimensions for the $k^{th}$ element (i.e., bin) as shown in Equation 5:
@@
In order to only keep “strong” edge responses, a global threshold, $\theta_{G}$, is selected and the voting function, $\mathcal{B}$, transforms the edge responses to “votes” for each orientation as shown in Equations 6 and 7. The first case to consider is when an edge response is to be recorded. The edge response must be larger than any other edge orientation and greater than the threshold. If both conditions are met, then a “1” is assigned to the corresponding edge bin as shown in Equation 6:
@@
Figure 4: Overall process for EHD feature. The image is first convolved with Sobel kernels [sobel19683x3] to generate edge responses for each orientation. The next step is to record the orientation that has the max response. Once the max response is recorded, a threshold is applied to ensure only “strong” edges are retained. The final step is to aggregate the max responses for each edge orientation to produce the final feature maps. The edge counts are binned into a histogram feature vector (similar to LBP).
@@
An example of EHD features is shown in Figure 4. “Engineered” histogram-based features such as LBP and EHD have the same limitations as traditional feature engineering approaches. These limitations include manually tuning parameters and domain expertise [nanni2017handcrafted]. To overcome these issues, others have proposed to 1) combine traditional and deep learning feature extraction approaches and 2) design neural networks to extract “engineered” features through network design.
@@
Figure 5: The proposed neural “engineered” feature pipeline for NEHD (Figure 5(a)) and NLBP (Figure 5(b)). The structural texture information is captured in the blue box and the statistical texture information is shown in the green box. In Figure 5(a), the input image is convolved with kernels to capture each edge orientation. The “no edge” operation can be implemented using a threshold or convolution layer. In Figure 5(b), the input image is convolved with kernels to capture the difference between neighboring pixels and the center pixel. The “threshold” operation is implemented using an activation function (e.g., sigmoid). After the “threshold” function, a $1\times 1$ convolution is used to compute a weighted sum operation of the difference maps to generate the “bit” map of LBP highlighted in orange. After the structural texture information is extracted, both features use a histogram layer to aggregate the structural texture information into statistical texture features.
II-B Combination of Traditional and Deep Learning Approaches
@@
Traditional feature extraction approaches and deep learning methods are used 1) separately and 2) together. Generally, there are five approaches for combining traditional and deep learning approaches [peeples2022connecting]: 1) take features from deep learning model and pass into traditional classifiers [scabini2019evaluating, sani2017learning, zhang2016svm], 2) “engineer” filters of CNN (may update through backpropagation or keep fixed) or emulate “engineered” features via the network design [bruna2013invariant, chan2015pcanet, malof2018improving, bianconi2019cnn, Su_2021_ICCV] , 3) pass “engineered” features into the network as input [muhammad2017tex, anwer2018binary, van2019feeding], 4) combine both “engineered” and CNN features for traditional classifiers [nguyen2018combining, paul2016combining, wu2016multi], and 5) texture encoding methods [cimpoi2015deep, zhang2017deep, song2017locally, xue2018deep, hu2019multi].
@@
There are several problems with existing approaches for combining traditional and deep learning methods. First, an issue with incorporating “engineered” features are that the models cannot be trained in an end-to-end fashion because the “engineered” features are not updated as performance changes. Also, there are additional computational costs for training deep learning model(s) and separate classifier(s). Along with the increased computational costs, more parameter tuning is necessary for the “engineered” features, deep learning model(s), and classifier(s). Lastly, deep learning features perform well in practice, but these features are not easily explainable and/or interpretable as traditional features though there are ongoing efforts to “open” the black box [rudin2022interpretable].
III Method
III-A Neural “Engineered” Features Histogram Layer
@@
The baseline histogram layer function in [peeples2021histogram] equally weights input features. As constructed, this implementation of the histogram layer will not be able to account for structural changes in texture information. For example, two binary images of a cross and checkerboard pattern will have the same distribution of pixels, but the arrangement of these pixels is different [peeples2022histogram]. The unweighted histogram layer will misidentify these two different structural texture types. To account for this, one can simply learn a weighting of the input features to account for structural differences. Our proposed neural “engineered” feature layer, $f$, takes an input image or feature map(s), $\mathbf{X}$, and applies two functions to extract texture information as shown in Equation 8:
@@
where $\phi$ and $\psi$ represents the statistical and structural texture information respectively. Generally, $\psi$ is implemented as a local feature extractor in a given neighborhood $\mathcal{N}$ (e.g., $3\times 3$ edge kernel) and $\phi$ is selected as a global or local operation.
$\phi$ represents the histogram layer introduced in [peeples2021histogram]. The histogram layer output is shown in Equation 9. The normalized frequency count, $\phi_{rcbk}$, is computed with a sliding window of size $S\times T$ and the binning operation for a histogram value in the $k^{th}$ channel of the input $x$:
@@
III-B Neural Edge Histogram Descriptor

### v2

@@ Abstract
In the computer vision literature, many effective histogram-based features have been developed. These “engineered” features include local binary patterns and edge histogram descriptors among others and they have been shown to be informative features for a variety of computer vision tasks. In this paper, we explore whether these features can be learned through histogram layers embedded in a neural network and, therefore, be leveraged within deep learning frameworks. By using histogram features, local statistics of the feature maps from the convolution neural networks can be used to better represent the data. We present neural versions of local binary pattern and edge histogram descriptors that jointly improve feature representation for image classification. Experiments are presented on benchmark and real-world datasets. Our code is publicly available  [^1] https://github.com/Advanced-Vision-and-Learning-Lab/NEHD_NLBP.
{IEEEImpStatement}
Our research introduces learnable histogram-based features in neural networks, such as neural local binary patterns (NLBP) and neural edge histogram descriptors (NEHD). These novel layers enable the integration of statistical image processing algorithms directly into deep learning frameworks. By leveraging these learnable histogram-based features within deeper networks, we can enhance feature representation capabilities, thereby improving performance in image classification tasks and potentially discovering new histogram-based features. Our publicly available code facilitates broader adoption and exploration of these techniques, contributing to advancements in computer vision applications that rely on robust, adaptive, and explainable feature extraction methods.
{IEEEkeywords}
@@
1 Introduction
Before the popularity of deep learning, feature engineering often (and still does) played a vital role in the fields of computer vision and machine learning. Examples of engineered features include local binary patterns [1] (LBP) and edge histogram descriptors (EHD) [2]. Extracting and selecting optimal features through feature engineering was often difficult. Many of these engineered features were created through the use of histograms, but there were many parameters (such as bin centers and widths) that were difficult to tune. As an alternative to the difficult and time-consuming process of feature engineering, deep learning is used to automate the process of extracting features and performing follow-on tasks (e.g., classification, segmentation, object detection).
Convolutional neural networks (CNN) are frequently used in a variety of applications. These models extract features through a series of convolutions, aggregation functions (e.g., max and average pooling) and non-linear activation functions to improve the representation of the data for better performance. Despite these powerful expressive features learned by a CNN, these models have some downfalls. CNNs have a spatial stationary assumption; therefore, CNNs are unable to account for changes in the local statistics (i.e., statistical texture features [3]) for regions of the data [4, 5]. Additionally, CNNs can lead to increased computational cost in terms of training time, memory requirements, and large datasets [6]. CNNs effectively capture structural texture features while histogram layers focus on statistical texture features [3].
To mitigate these issues associated with traditional and deep learning features, alternative models have been introduced that take inspiration from both approaches. A notable method is the local binary convolutional neural network (LBCNN) [6]. The LBCNN used a novel architecture design that led to a less expensive model that performed comparably to standard CNNs. LBCNN was inspired by LBP; however, the LBCNN generalizes the code generation and does not account for the aggregation operation of the original LBP.
Figure 1: Feature maps captured by LBP (top) and NLBP (bottom) from the FashionMNIST dataset. The proposed approach nearly reconstructs the original LBP feature. The LBP and NLBP encodings differ slightly due to the non-linearity: threshold (LBP) and ReLU (NLBP). Additionally, a square window ($\times 3$ 3) is used for NLBP while LBP used a radial window for the local neighborhood. There are small magnitude differences of the histogram because of the soft binning approximation used in NLBP.
@@ 1 Introduction
Histogram-based feature approaches (e.g., LBP and EHD) as well as CNNs are able to extract texture information [7, 8, 9, 10] to represent the data. There are two types of textures: statistical and structural [10, 11]. Both texture feature types are useful to represent the data to improve performance and features such as LBP use both within the feature extraction pipeline [7]. In this work, we present a generalized framework to learn histogram-based features in artificial neural networks: neural LBP (NLBP) and EHD (NEHD). The contributions of this new method are four-fold:
@@
Each neural feature can be reconstructed using convolutional and histogram layers to learn the structural and statistical texture features respectively. Our proposed neural “engineered” features can closely reconstruct the LBP and EHD features as shown in Figures 1 and 2. The LBP feature consists of both structural (i.e., encoded pixel differences) and statistical (i.e., binary code frequency) texture information. To capture the structural relationships, following [12], sparse convolutional kernels can be used to capture the difference between the center and neighboring pixels in an image. After the convolution operation is performed, an activation function (e.g., sigmoid, ReLU) can be used instead of the threshold operation. After the input image has been encoded, a histogram layer can be used to adaptively aggregate the encoded features as opposed to the fixed binning of the original LBP feature.
Figure 1 shows a comparison of LBP and the proposed NLBP. To reconstruct the LBP feature, a modified ReLU activation function can be used. The standard ReLU function sets all negative values to 0 (similar to the threshold function in LBP), but if the center pixel and neighbor pixel value are the same, the LBP threshold function will output a value of 111. To account for this in the NLBP, if the difference between a center and neighboring pixel is 0, the ReLU function was set to output a 111 instead of a 0 resulting in a similar encoding map as the original LBP. The histogram layer bins are initialized to the corresponding LBP code values ranging from 0 to 255255255. The bin width was set to 3.753.753.75 as this value will correspond to a narrow histogram bin.
The EHD feature also captures structural (i.e., edges) and statistical (i.e., max response frequency) texture information. The structural features can be captured using a convolutional layer with each filter corresponding to a Sobel kernel for each edge orientation, $0^{\circ}$ to $315^{\circ}$. A total of eight edge kernels were used to capture horizontal, vertical, diagonal and anti-diagonal orientations (as in the standard EHD feature). To account for direction information, the bin center values are set to the maximum response of the convolution output from the edge kernels and the input data (i.e., maximum input value $\times$ ($+1$ 2 + 1)). By changing the bin center, the binning function will now account for direction information. The anti-direction will now have the lowest response to the corresponding correct direction (i.e., $0^{\circ}$ will have lowest response for $180^{\circ}$). The same bin width value (3.753.753.75) was used for NEHD as in NLBP. The parameters (convolution kernels, bin centers, and bin widths) were fixed to ensure that the histogram layer can mimic the EHD feature as shown in Figure 2.
@@
2 Related Work
Figure 3: Example of LBP feature value calculation. After a local neighborhood is selected (e.g., $\times 3$ 3), the difference between the center and neighboring pixels is computed. If the difference is greater than or equal to a chosen threshold (e.g., 0), then a value of “1” is returned for that neighbor. The next step is to encode the binary position of each neighbor (i.e. $2^{0}$, $2^{1}$, … , $2^{7}$). An elementwise multiplication and sum operation is performed between the threshold and binary position values. The output of this operation is the LBP feature for the local patch. The values from each local patch of the input image is then aggregated via a histogram to generate the output LBP histogram vector.
2.1 “Engineered” Histogram Features
Histograms are used throughout computer vision and machine learning as a method to aggregate intensity and/or feature values as well as relationships between neighboring inputs (e.g., edge orientation, pixel differences). Common histogram-based features include LBP [1], histogram of oriented gradients (HOG) [13], EHD [14], and Haralick texture features extracted from the gray level co-occurrence matrix (GLCM) [15]. For this work, we will focus on LBP and EHD.
@@ 2 Related Work
LBP is computed by first selecting a neighborhood of size $\mathcal{N}$ (commonly 3 $\times$ 3 or also using a radial window operation) and computing the difference between a center pixel, $x_{c}$, and each of the pixels in the defined neighborhood, $x_{i}$, in a gray scale image. The LBP feature value for a pixel in an image is computed by computing the product of the threshold, $\mathcal{T}$, of the difference between the center pixel and neighbor pixel, and two raised to the power of the index of the neighbor (Equation 1). The threshold for assigning a “0” or a “1” is usually 0 (i.e., the difference needs to be positive) as shown in Equation 2, but other variants may use different thresholds [16, 17]:
@@
LBP is robust to monotonic changes in grayscale due to the threshold function [1, 19, 20] (i.e., if the intensity values of an image changes, the binary threshold will always return either two values, “0” or “1”, despite gray level changes). There are a plethora of extensions to LBP in the literature for different applications including understanding cell phenotypes leveraging microscopy images [21] and for biomedical applications such as pap smear classification [22]. Several works are available that summarize the novelty of different LBP-based approaches [12, 23]. Work continues to be done to make improvements for the LBP feature in particular and our method aims to contribute to ongoing work related to this feature.
@@
The EHD feature records the frequency and orientation of intensity changes in an image [14]. A single channel input image is convolved with a set of $K$ filters (e.g., Sobel [24]) of size $M\times N$ to calculate the edge responses, $R\in\mathbb{R}^{M^{\prime}\times N^{\prime}\times K}$. The edge responses are grouped (“binned”) together based on their orientation such as vertical, horizontal, diagonal, anti-diagonal, and no edges (isotropic) [14]. The five groups can include signed [25, 26] or unsigned orientations [14] (e.g., 0 and 180 are both horizontal but different directions). To compute the histogram of edge orientations, $\mathbf{H}\in\mathbb{R}^{K+1}$, the maximum edge responses are summed over the spatial dimensions for the $k^{th}$ element (i.e., bin) as shown in Equation 5:
@@
In order to only keep “strong” edge responses, a global threshold, $\theta_{G}$, is selected and the voting function, $\mathcal{B}$, transforms the edge responses to “votes” for each orientation as shown in Equations 6 and 7. The first case to consider is when an edge response is to be recorded. The edge response must be larger than any other edge orientation and greater than or equal to the threshold. If both conditions are met, then a “1” is assigned to the corresponding edge bin as shown in Equation 6:
@@
Figure 4: Overall process for EHD feature. The image is first convolved with Sobel kernels [24] to generate edge responses for each orientation. The next step is to record the orientation that has the max response. Once the max response is recorded, a threshold is applied to ensure only “strong” edges are retained. The final step is to aggregate the max responses for each edge orientation to produce the final feature maps. The edge counts are binned into a histogram feature vector (similar to LBP).
@@
An example of EHD features is shown in Figure 4. “Engineered” histogram-based features such as LBP and EHD have the same limitations as traditional feature engineering approaches. These limitations include manually tuning parameters and domain expertise [27]. To overcome these issues, others have proposed to 1) combine traditional and deep learning feature extraction approaches and 2) design neural networks to extract “engineered” features through network design.
@@
Figure 5: The proposed neural “engineered” feature pipeline for NEHD (Figure 5(a)) and NLBP (Figure 5(b)). The structural texture information is captured in the blue box and the statistical texture information is shown in the green box. In Figure 5(a), the input image is convolved with kernels to capture each edge orientation. The “no edge” operation can be implemented using a threshold or convolution layer. In Figure 5(b), the input image is convolved with kernels to capture the difference between neighboring pixels and the center pixel. The “threshold” operation is implemented using an activation function (e.g., sigmoid). After the “threshold” function, a $\times 1$ 1 convolution is used to compute a weighted sum operation of the difference maps to generate the “bit” map of LBP highlighted in orange. After the structural texture information is extracted, both features use a histogram layer to aggregate the structural texture information into statistical texture features.
2.2 Combination of Traditional and Deep Learning Approaches
@@
Traditional feature extraction approaches and deep learning methods are used 1) separately and 2) together. Generally, there are five approaches for combining traditional and deep learning approaches [28]: 1) take features from deep learning model and pass into traditional classifiers [29, 30, 31], 2) “engineer” filters of CNN (may update through backpropagation or keep fixed) or emulate “engineered” features via the network design [32, 33, 34, 35, 36] , 3) pass “engineered” features into the network as input [37, 38, 39], 4) combine both “engineered” and CNN features for traditional classifiers [40, 41, 42], and 5) texture encoding methods [43, 44, 45, 46, 47].
@@
There are several problems with existing approaches for combining traditional and deep learning methods. First, an issue with incorporating “engineered” features are that the models cannot be trained in an end-to-end fashion because the “engineered” features are not updated as performance changes. Also, there are additional computational costs for training deep learning model(s) and separate classifier(s). Along with the increased computational costs, more parameter tuning is necessary for the “engineered” features, deep learning model(s), and classifier(s). Lastly, deep learning features perform well in practice, but these features are not easily explainable and/or interpretable as traditional features though there are ongoing efforts to “open” the black box [48].
3 Method
3.1 Neural “Engineered” Features Histogram Layer
@@
The baseline histogram layer function in [10] equally weights input features. As constructed, this implementation of the histogram layer will not be able to account for structural changes in texture information. For example, two binary images of a cross and checkerboard pattern will have the same distribution of pixels, but the arrangement of these pixels is different [3]. The unweighted histogram layer will misidentify these two different structural texture types. To account for this, one can simply learn a weighting of the input features to account for structural differences. Our proposed neural “engineered” feature layer, $f$, takes an input image or feature map(s), $\mathbf{X}$, and applies two functions to extract texture information as shown in Equation 8:
@@
where $\phi$ and $\psi$ represent the statistical and structural texture information, respectively. Generally, $\psi$ is implemented as a local feature extractor in a given neighborhood $\mathcal{N}$ (e.g., $\times 3$ 3 edge kernel) and $\phi$ is selected as a global or local operation.
$\phi$ represents the histogram layer introduced in [10]. The histogram layer output is shown in Equation 9. The normalized frequency count, $\phi_{rcbk}$, is computed with a sliding window of size $S\times T$ and the binning operation for a histogram value in the $k^{th}$ channel of the input $x$:
@@
3.2 Neural Edge Histogram Descriptor